OpenStudy (anonymous):

in order to achieve maximum results from aerobic exercise one should maintain ones heart rate at a certain level a 40 yr old woman with a resting of 80 beats per miniute should keep her heart rate between 140 and 160 beats per minute while exercising she checks her pulse at 10 seconds while exercising what should the beats be the number of beats should be between blank and blank beats

OpenStudy (anonymous):

140/6 and 160/6

OpenStudy (anonymous):

thats not right it says if her heart rate should be between 140 and 160 beats per min what should the lower and upper numbers be for 10 seconds?

OpenStudy (anonymous):

Yeah. Divide those by 6.

OpenStudy (anonymous):

\[140 \le X \le 160\]60/10=6\[\frac{140}{6}=23.3\]\[\frac{160}{6}=26.7\]\[23.3 \le x \le 26.7\]

OpenStudy (anonymous):

so 23 and 26 ?

OpenStudy (anonymous):

round 27 23 and 27

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

heres one similar a 35 year old woman with a resting heart rate of 80 beats per min should keep her heart rate between 143 and 164 beats per min while exercising

OpenStudy (anonymous):

she clocks her pulse for 10 sec while exercising what should the number of beats be ? the num of beats in 10 sec should be between blank and blank hint round to the nearest whole number

OpenStudy (anonymous):

Divide 143 and 164 by 6 again.

OpenStudy (anonymous):

\[\frac{x}{10}=\frac{143}{60}\]\[60x=143*10\]\[x=\frac{143}{10}=23.83\] \[\frac{x}{10}=\frac{164}{60}\]\[60x=164*10\]\[x=27.33\]round\[24 \le x \le 27\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

urwc, if u like my answer, click good answer,ty

OpenStudy (anonymous):

where at do i click that ?

OpenStudy (anonymous):

can u help me with another one please i have been stuck on them all night

OpenStudy (anonymous):

THE OTHER ONE IS BAY SIDE INSURANCE OFFERS TWO HEALTH PLANS , UNDER PLAN A GISELLE WOULD HAVE TO PAY THE FIRST $190 OF HER MEDICAL BILLS PLUS 40% OF THE REST UNDER PLAN B GISELLE WOULD PAY THE FIRST 250BUT ONLY 25 % OF THE REST FOE WHAT AMOUNT OF MEDICAL BILLS WILL PLAN B SAVE GISELLE MONEY? ASSUME SHE HAS 250 IN BILLS SO IT WOULD BE GISELLE WOULD SAVE WITH PLAN B IF SHE HAD MORE THAN BLANK IN BILLS ROUND TO THE NEAREST WHOLE NUMBER